OpenStudy (anonymous):

Rewrite the following equation into logarithmic form. 3^(x+2)=81

OpenStudy (anonymous):

\[3^{x+2}=81\] \[\log_3(3^{x+2})=\log_381\] \[(x+2).\log_33=\log_381\] \[(x+2)=4\] \[x=4-2\] \[x=2\]

OpenStudy (anonymous):

okay great thats what i got too. (: